# LoRA is All You Need for Safety Alignment of Reasoning LLMs

### v3

@@ 7 Conclusion
In this work, we identified a simple yet effective solution to the previously observed tension between reasoning and safety in LLMs. Through extensive experiments, we showed that applying LoRA during safety fine-tuning preserves reasoning capabilities while achieving strong safety alignment, in contrast to full-model fine-tuning, which significantly degrades reasoning. We further ablate the LoRA configuration to identify the key elements and make three important findings: (1) rank-$1$ updates are sufficient to achieve the best reasoning–safety tradeoff, (2) the up projection modules are most important for the tradeoff, yielding even better results than updating all or other modules, and (3) middle layers matter most, more than early or late layers. These results highlight that strong safety and reasoning can be achieved at minimal computational cost when updates are applied in the right places. Finally, we analyze the weight structure of LoRA and find that its updates are more orthogonal to the initial weights, which may help explain their reduced interference with reasoning. While our attempts to further enforce orthogonality yield only modest and inconsistent improvements, they highlight the potential of better controlling the geometry of safety updates as an avenue for future research. Overall, our results provide both a practical recipe for safety alignment of reasoning models and new insights into why LoRA is uniquely well-suited to this problem.
@@ A.1 Details about DirectRefusal
A.2 Details about StrongReject
@@ A.3 Details about AIME and GPQA
We adapted the codebase from Li et al. (2025b), which builds on top of lm-evaluation-harness Gao et al. (2024), to evaluate models on GPQA and AIME. For AIME, answers are evaluated using Qwen2.5-32B-Instruct as a judge. For GPQA, since it is a multiple-choice benchmark, answers are evaluated via regular expression matching. To improve evaluation accuracy, we made some small adjustments. One change is modifying the prompt to better ensure the model outputs answers in the desired format:
@@ A.4 Details about HumanEval and MBPP
We adapted the codebase of EvalPlus Liu et al. (2023). The original implementation included a response prefix, designed for earlier models that did not explicitly support an intermediate reasoning process. This prefix—such as “Below is a Python script with a self-contained function that efficiently solves the problem and passes corresponding tests:”—was prepended to model outputs during generation. However, this practice introduces unfair bias by encouraging all models to directly generate code. It inadvertently benefits the full-model fine-tuned baseline—which would otherwise often refuse to answer—by effectively forcing it to produce code. Conversely, it disadvantages reasoning-aligned models by disrupting the expected format that includes intermediate “thought”, causing them to skip the thinking process entirely. This can result in skewed conclusions. We suspect this explains the abnormally low accuracy of the base reasoning model and the high accuracy of the full-model fine-tuned variant reported in Jiang et al. (2025). To address this issue, we remove the response prefix. We also slightly reword the prompt to better align with reasoning models.
@@ A.5 Training Details
For 7B models, full-model fine-tuning is conducted using 4 GPUs with a batch size of 2 per device for 5 epochs. LoRA fine-tuning uses 2 GPUs with a batch size of 2 per device for 10 epochs. We set the LoRA hyperparameters as $\alpha=16$ and lora_dropout = 0.05.
For 14B models, full-model fine-tuning is performed using 8 GPUs with a batch size of 1 per device for 5 epochs. LoRA fine-tuning uses 4 GPUs with a batch size of 2 per device for 10 epochs. We set the LoRA hyperparameters as $\alpha=16$ and lora_dropout = 0.05.
@@
Appendix B Detailed Description of the Methods Used in Section 6.2
We consider two approaches for enforcing stronger orthogonality: regularization during LoRA fine-tuning and enforcing orthogonality during merging LoRA weights, as we discuss next.
Regularization during LoRA training. We add a penalty term to the loss that discourages overlap between ${\bm{W}}_{I}$  [^1] To avoid out-of-memory issues during training caused by the large dimensionality of model weights, in implementation we use a low-rank approximation of ${\bm{W}}_{I}$ instead of the full matrix. and $\Delta{\bm{W}}$, specifically:
  • Regularization-col: $\beta(\frac{\|{\bm{W}}_{I}^{\top}\Delta{\bm{W}}\|}{\|{\bm{W}}_{I}\|\|\Delta{\bm{W}}\|)})^{2}$ encourages orthogonality in the column space.
  • Regularization-both: $\beta(\frac{\|{\bm{W}}_{I}^{\top}\Delta{\bm{W}}\|}{\|{\bm{W}}_{I}\|\|\Delta{\bm{W}}\|)})^{2}+\beta(\frac{\|\Delta{\bm{W}}^{\top}{\bm{W}}_{I}\|}{\|{\bm{W}}_{I}\|\|\Delta{\bm{W}}\|})^{2}$ encourages orthogonality in both column and row spaces.
@@
Enforcing Orthogonality During Merging LoRA Weights. Starting with updates obtained from standard LoRA, we modify how they are merged with ${\bm{W}}_{I}$, which we call OrthoMerge. Before applying Equation 1, we preprocess $\Delta{\bm{W}}$ as follows:

### v4

@@ 7 Conclusion
Applying LoRA during safety alignment of reasoning LLMs can effectively mitigate the previously observed “Safety Tax”, achieving strong safety without compromising reasoning, as validated by our extensive experiments. We further ablate LoRA configurations and find that (1) rank-$1$ updates are sufficient and often optimal, (2) the MLP up-projection layers are the most important, and (3) middle layers matter more than early or late layers, providing practical guidance for achieving strong performance at minimal fine-tuning cost. Finally, our theoretical analysis clarifies when and why LoRA works, revealing novel insights into how rank overshoot interacts with the intrinsic dimensionality of the base capability.
Acknowledgement
This research was supported in part by in part by the NSF CAREER Award 2146492, NSF-Simons AI Institute for Cosmic Origins (CosmicAI), and NSF AI Institute for Foundations of Machine Learning (IFML).
@@ A.1 Details about DirectRefusal
A.2 Details about Safety Evaluation
@@ A.3 Details about AIME and GPQA
We adapted the codebase from (Li et al., 2025b), which builds on top of lm-evaluation-harness (Gao et al., 2024), to evaluate models on GPQA and AIME. For AIME, answers are evaluated using Qwen2.5-32B-Instruct as a judge. For GPQA, since it is a multiple-choice benchmark, answers are evaluated via regular expression matching. To improve evaluation accuracy, we made some small adjustments. One change is modifying the prompt to better ensure the model outputs answers in the desired format:
@@ A.4 Details about HumanEval and MBPP
We adapted the codebase of EvalPlus (Liu et al., 2023). The original implementation included a response prefix, designed for earlier models that did not explicitly support an intermediate reasoning process. This prefix—such as “Below is a Python script with a self-contained function that efficiently solves the problem and passes corresponding tests:”—was prepended to model outputs during generation. However, this practice introduces unfair bias by encouraging all models to directly generate code. It inadvertently benefits the full-model fine-tuned baseline—which would otherwise often refuse to answer—by effectively forcing it to produce code. Conversely, it disadvantages reasoning-aligned models by disrupting the expected format that includes intermediate “thought”, causing them to skip the thinking process entirely. This can result in skewed conclusions. We suspect this explains the abnormally low accuracy of the base reasoning model and the high accuracy of the full-model fine-tuned variant reported in (Jiang et al., 2025). To address this issue, we remove the response prefix. We also slightly reword the prompt to better align with reasoning models.
@@ A.5 Training Details
For 7B models (Qwen) and 8B models (Llama), full-model finetuning is conducted using 4 GPUs with a batch size of 2 per device for 5 epochs. LoRA finetuning uses 2 GPUs with a batch size of 2 per device for 10 epochs. We set the LoRA hyperparameters as $\alpha=16$ and lora_dropout = 0.05.
For 14B models, full-model finetuning is performed using 8 GPUs with a batch size of 1 per device for 5 epochs. LoRA finetuning uses 4 GPUs with a batch size of 2 per device for 10 epochs. We set the LoRA hyperparameters as $\alpha=16$ and lora_dropout = 0.05.
@@
Appendix B Additional Plots for Experiments in Section 4
B.1 Results on BeaverTails
The results on BeaverTails across models are shown in Figure 11, and the observed pattern is consistent with that in Figure 3.
(a) R1-Distill-Qwen-7B
(b) R1-Distill-Qwen-14B
(c) R1-Distill-Llama-8B
Figure 11: Results with safety measured on BeaverTails, consistent with the pattern shown in Figure 3.
B.2 Results on the Base Versions of HumanEval and MBPP
Figure 13 shows the results on the base versions of HumanEval and MBPP.
(a) R1-Distill-Qwen-7B
(b) R1-Distill-Qwen-14B
(c) R1-Distill-Llama-8B
Appendix C Additional Experimental Results: Exploring the Structure of LoRA Weights
We empirically study the structure of LoRA weight updates. Preliminary results suggest that LoRA induces updates with slightly smaller overlap with the initial weights than full-model finetuning, which may indicate that LoRA is less disruptive to reasoning-related weights. We then explore methods to explicitly reduce this overlap (via regularization or post-hoc weight merging). One post-hoc method yields a modest improvement in the reasoning–safety trade-off on some tasks. These results are preliminary, and further work is needed to develop approaches that produce consistent gains across tasks; we hope our findings will inspire future work in this direction.
(a) $\frac{\|{\bm{W}}_{I}^{\top}\Delta{\bm{W}}\|}{\|{\bm{W}}_{I}\|\|\Delta{\bm{W}}\|}$
(b) $\frac{\|{\bm{U}}_{16}{\bm{U}}_{16}^{\top}\Delta{\bm{W}}\|}{\|\Delta{\bm{W}}\|}$
(c) $\frac{\|{\bm{W}}_{I}\Delta{\bm{W}}^{\top}\|}{\|{\bm{W}}_{I}\|\|\Delta{\bm{W}}\|}$
Figure 14: LoRA updates exhibit smaller overlap with the original weights compared to the full-model finetuning updates. Although the reduction in overlap is sometimes small, it can be observed across most layers for all four metrics, which cover both the column (a)(b) and row spaces (c)(d). The 14B models are used here.
C.1 LoRA Updates Have Less Alignment with Initial Weights
We compare the full-rank fine-tuned model with the LoRA fine-tuned model in which both the attention and MLP modules are updated with $r=4$, making the two settings more comparable since updates occur in all major modules. We compute the alignment metrics for different module types across layers, average them over layers, and report the results for the 14B models in Figure 14. We observe that LoRA achieves smaller overlap in most modules across the metrics, with a few exceptions. This suggests that, for most weights, ${\bm{W}}_{I}$ and $\Delta{\bm{W}}$ are more orthogonal—both in the column and row spaces—for LoRA than for full-model finetuning. In other words, under LoRA finetuning, the safety-oriented updates read from and write into subspaces that are more separate from those used by the original reasoning-related weights—more so than in the full-model fine-tuned version. Although the reduction in alignment values is sometimes small, it may still indicate that LoRA updates interfere less with the reasoning-related components of the model, potentially explaining the better preservation of reasoning performance. A more in-depth investigation is needed to fully understand the underlying mechanisms and to develop more precise metrics for capturing this effect—an important direction for future work.
(a) 14B models on AIME and GPQA.
(b) 14B models on HumanEval+ and MBPP+.
C.2 Exploring Methods That Further Reduce Alignment
Given the observations in Section C.1, we ask whether further reducing the overlap between $\Delta{\bm{W}}$ and ${\bm{W}}_{I}$ could lead to even better reasoning performance without compromising safety. This question is particularly relevant because, while the results achieved by LoRA in Section 4.4 are strong, they are not perfect—a small performance gap remains compared to the original reasoning model, especially for the 14B model on AIME, HumanEval+, and MBPP+.
We experiment with two approaches:
  • Regularization during LoRA training. We add a penalty term to the loss that discourages overlap between ${\bm{W}}_{I}$ and $\Delta{\bm{W}}$ (to avoid out-of-memory issues during training caused by the large dimensionality of model weights, in implementation we use a low-rank approximation of ${\bm{W}}_{I}$ instead of the full matrix), specifically:
    • reg-col: $\beta(\frac{\|{\bm{W}}_{I}^{\top}\Delta{\bm{W}}\|}{\|{\bm{W}}_{I}\|\|\Delta{\bm{W}}\|)})^{2}$ encourages orthogonality in the column space.
    • reg-both: $\beta(\frac{\|{\bm{W}}_{I}^{\top}\Delta{\bm{W}}\|}{\|{\bm{W}}_{I}\|\|\Delta{\bm{W}}\|)})^{2}+\beta(\frac{\|\Delta{\bm{W}}^{\top}{\bm{W}}_{I}\|}{\|{\bm{W}}_{I}\|\|\Delta{\bm{W}}\|})^{2}$ encourages orthogonality in both column and row spaces.
@@ C.2 Exploring Methods That Further Reduce Alignment
  • Post-hoc Orthogonalization: We post-process $\Delta{\bm{W}}$ before applying Equation 1, projecting it onto the orthogonal complement of the top-$k$ singular vectors of ${\bm{W}}_{I}$: